a spherical watermelon of radius 4cm is halved. the uniform width of the white part surrounding the red edible part is 1cm. from half of the melon, the red part of the fruit is scooped out. find (i)the volume of the edible part of the fruit taken out (ignore the volume of seeds). take pi=3.14
(ii) the surface area of the white part of that hemispherical piece

Answer by mananth(16949)


Volume of hemisphere = (2/3) pi*r^3
here r=3
Volume of edible part= 2/3 * pi * 3^3= 18 pi
18*3.14 cm^3
Surface area of white part = 2 *pi*r^2
2*pi*3^2
= 18 pi cm^2
18*3.14 cm^2